# Distance-Preserving Spatial Representations in Genomic Data

### v2

@@ Abstract
Representation learning, spatial transcriptomics, single-cell RNA sequence, variational inference, and dimensionality reduction.
Figure 1: The model architecture of dp-VAE with vanilla VAE as its backbone. During the training stage, the distance-preserving loss ($\mathcal{L}_{\textrm{DP}}$) enforces distance preservation between the spatial domain ($S$) and the latent space ($Z$), enforcing the encoder network (e.g., parameterized as a fully connected neural network) to capture meaningful signals of the tissue spatial characteristics of the provided gene expression ($X$). During the inference/test stage, the extracted latent representations are fed to an optimizer, such as multidimensional scaling (MDS) if it is a spatial reconstruction task instead of a spatial context imputation task.
@@ 1 Introduction
Understanding the tissue spatial context from which the gene expressions are extracted is crucial to many applications, deepening the understanding of cellular organizations and their interactions within tissues, as well as illuminating biological insights in neuroscience, developmental biology, precise diagnostics, personalized therapies, and a range of disease mechanisms such as cancer [57]. As a result, measurement technologies such as spatial transcriptomics have emerged as a powerful tool in modern biology, bridging the gap between single-cell genomics and the larger physiological processes at play in living tissues.
However, various practical and technical constraints limit the range of spatial context and the granularity that these measurement tools can directly capture. For example, spatial transcriptomics (ST) technologies such as Visium by 10x Genomics and NanoString GeoMx [10] typically do not have single-cell resolution [7], making the extracted spatial coordinates within the tissue space too coarse for many precision-focused downstream analyses. Enhancing resolution to this degree requires specialized equipment, intricate protocols, and substantial computational infrastructure, thereby limiting its availability to highly specialized institutions and facilities. On the other hand, the usual single-cell RNA sequencing (scRNA-seq) is capable of extracting gene expression at a localized single-cell level but lacks spatial contexts. Given observations of the complementary nature of the pro-and-cons of these two existing mainstream gene expression extraction technologies, the following question can be asked:
@@
An approach towards solving this question can potentially avoid the need for the deployment of spatial transcriptomic, making the tasks we intend to carry out with spatial context to be more accessible.
As an initial step toward tackling this question, we propose a generic representation learning and transfer learning framework enabling the inference of spatial context from purely gene expression data (e.g., collected from scRNA-seq) by drawing reference from datasets where gene expression is paired with spatial information (e.g., collected from spatial transcriptomics). This procedure requires little or even no reliance on ST technologies depending on the type of reference data that the user intends to use, which can even come from existing publicly available databases. Specifically, our representation learning framework is based on the variational autoencoder (VAE) models [23], consisting of an encoder and a decoder network and trained on the reference datasets. Instead of using the standard ELBO loss, we propose to regularize it with a distance-preserving loss function, which utilizes the spatial context as auxiliary information to enforce the learned representation to be geometrically similar to the reference dataset. This regularization helps the model capture useful marker genes and signals from gene expressions that are representative of their spatial contexts. During inference, since spatial context is not used as the input to the model, the model can be applied to gene expression datasets without spatial context information but may share similar expression and geometric patterns with the reference dataset, thereby effectively extrapolating to the task of spatial reconstruction and imputation of the missing spatial context in these inference datasets. Next, the spatial reconstruction task can be formulated as an optimization that reversely solves for the optimal coordinates sets aligning with the distance geometry of extracted latent representations from the input gene expression data. This optimization can be bypassed when we set the latent representation space to match the dimensionality of the tissue’s spatial domain. An illustration of the pipeline is provided in Figure 1.
@@ 2 Related Works
Spatial transcriptomics analysis has consistently garnered interest from the statistical machine learning community. It includes different tasks such as cellular deconvolution, clustering, visualization, cell classification, batch correction, gene imputation, and differential expression [35, 31, 33, 13]. Classical statistical models such as marked point process [14], Gaussian process [46], generalized linear spatial models [45, 61], and autoregression [35] have been deployed for these modeling tasks, where spatial context information are typically integrated into these methods in the form of covariance matrices. In the meantime, machine learning models empowered with deep neural networks can instead directly leverage the spatial context information. For example, convolution neural network-based methods such as CoSTA [54], and/or graph-based methods using graph neural networks such as GLISS [62], SpaGCN [20], GraphST [31]. See [58] for a survey of these methods.
Our work aligns more closely with literature that explores how various generative models (e.g. autoencoders) can be applied to learning efficient representation from ST and scRNA-seq, such as STAGATE [9], scVSC [50], scVI [33] and gimVI [32]. Some studies have also proposed spatial reconstruction methodologies of tissue spatial structure, such as Tangram [5], and NovoSpaRc [38]. However, to the best of our knowledge, our work is the first to propose to bridge representation learning with tissue spatial reconstruction drawing reference from out-of-sample and out-of-distribution datasets. Additionally, the previously mentioned statistical and machine learning models dedicated to spatial transcriptomics are also inapplicable to such tasks, as the spatial contexts are required as a conditional input to the models during the inference stage.
The distance-preserving loss used in our study has (concurrently) emerged in parallel domains within the broader machine-learning literature. For instance, distance-preserving mappings have been introduced for GANs for unsupervised domain mapping [2], image-to-audio transformation [40, 22]. It has also appeared as regularizers for VAEs [6] and adversarial learning [18, 37]. A stronger notation compared to distance-preserving, named isometry [56, 26, 4] has also been extensively studied in machine learning models such as VAEs. Building on top of these works, our derivation of distance-preserving loss’ connection to distortion or the bi-Lipschitz condition contributes to formalizing this notion.
@@ 3 Problem setup
At the high level, the goal of representation learning is to learn a mapping that extracts latent embeddings $z\in\mathbb{R}^{d}$ from given gene expression $x\in\mathbb{R}^{D}$. In our setting, we additionally assume that we have access to some spatial context information $s\in\mathbb{R}^{2}$ associated with each gene expression indicating where the cell is located in the tissue slice. It is only available in the training stage, but not when during inference. We expect that the learned representation learning model can also undergo some procedure to reconstruct the spatial domain given only gene expression during the inference stage. Formally, given the gene expression a set of gene expressions denoted as $x_{i}\in\mathbb{R}^{D}$, $i=1,\ldots,N$, the goal of spatial reconstruction is to obtain a mapping to infer the spatial coordinates $s_{i}\in\mathbb{R}^{2}$ for each cellular location $i$.
In some cases, this task can be framed as an imputation task, where the spatial information of some cell has already been observed, indexed by $i\in\mathcal{I}\subseteq\left\{1,\ldots,N\right\}$. Therefore, imputation aims to infer the rest of the spatial coordinates locations $s_{i}$, where $i\in\left\{1,\ldots,N\right\}/\mathcal{I}$. We remark that, in practice, the spatial reconstruction tasks here are useful for many downstream analyses, such as spatial domain identification, region-region interaction analysis, and spatial trajectory analysis. The inference task is similar in spirit to the generic task of transfer learning [39].
@@ 4 Method
To tackle the aforementioned challenges, we propose a gene expression representation learning framework. The model, named distance-preserving variation autoencoder (dp-VAE), incorporates the spatial information into the learning objective as a penalty for the error of the pairwise similarity between the model’s learned latent variable with the spatial information of the training data. In this case, the spatial information can also be viewed as an auxiliary variable during the training stage rather than a direct input of the model, therefore it is not required during the inference stage [^1] Whereas conditional models such as CVAE [43, 52] do require the spatial context as inputs, therefore cannot be applied in this setting.. During inference, spatial reconstruction, and spatial imputation, this penalty can be used reversely to decipher the spatial coordinates from the extracted latent representations of the data.
@@ 4.1 Background: Variational Autoencoder
However, the existing encoding paradigm presented in (2) reveals implicit biases, undermining its efficacy. Two primary concerns highlight this limitation: ($i$) The likelihood-focused optimization objective, while capturing some meaningful signals, underscores probabilistic aspects that could be incorrectly specified due to cross-dependencies or non-identical distributed genomic samples. This could lead to entangled and biased representations of spatial context to be learned from the training data; ($ii$) In the meantime, directly inserting the spatial information as part of the input variable violates our assumption about the setting where spatial context is not available as input during the inference stage, but rather should be the output of the model. Therefore, there is a need to develop a new method that incorporates likelihood-based generative modeling strategies with spatial information.
@@ 4.2 Distance-Preserving Variational Autoencoder
where the reconstruction term $\mathcal{L}_{\rm recon}$ and KL divergence term $\mathcal{L}_{\rm KL}$ are as defined in (2). The outer-level regularization coefficients $\alpha_{1},\alpha_{2}\leq 0$ control for the numerical balance of the distance-preserving loss and the ELBO loss, and $\beta$ is the temperature coefficient for the KL-divergence term [19]. The model is trained by learning the parameters of the encoder and decoder networks $\theta$ and $\phi$ as well as the scaling parameter $\lambda$ to minimize (4).
@@
  • ($i$)
    From the biological perspective, neighboring cells, often from similar tissue types, tend to exhibit homogeneity in attributes like gene expression and biological functions. Representation-based generative models, which aim to capture such spatial relationships and tissue-level patterns, should utilize such geometric structures in learning disentangled representations that are meaningful. Prior works [20, 45] have illustrated that proximity in physical distance between cells has a strong association with their gene expressivity. This is also the initial of many spatial transcriptomics machine learning models to incorporate spatial context as inputs to enhance the modeling of cellular behavior and tissue organization, enabling more robust, context-aware insights into biological systems. If viewed reversely, it is also reasonable to claim that the learned gene expression representations can also reveal the associated spatial context from where they are located in the tissue.
@@
It is noted that the first point is also supported by a simple experiment we conduct, illustrated in Figure 2. It can be seen that out of the 27 datasets that we are using (see the experiment section for a detailed description of them), 23 datasets show a statistically significant ($95\%$) positive correlation between the spatial distance of the sample locations and the gene expression similarity. This highlights the potential of seeing the spatial context as a useful source of information to aid the learning of gene expression representations, or reversely, gene expression representations are capable of revealing information about their associated spatial context.
@@
~$\widetilde{\mathcal{L}}_{\rm DP}=\frac{1}{n^{2}}\sum_{i,j}\mathbf{G}_{i,j}\left|\|z_{i}-z_{j}\|-\lambda\|s_{i}-s_{j}\|\right|,$ (5)
@@ 4.3 Spatial Reconstruction and Imputation
In the next step, we aim to carry out the spatial reconstruction and imputation task, where we only have inference data consisting of gene expression but not spatial context information. Assuming that we have access to some gene expression data that are paired with spatial context information and are statistically similar to the inference data, then dp-VAE can be trained on reference dataset, but transfered to the inference data for application. Specifically, since dp-VAE encodes the spatial context patterns in the latent representation via a distance-preserving regulated mapping, we can reversely find the optimal spatial coordinates in the spatial domain that align with the distance patterns in the extracted latent space. This can be formulated as an optimization problem regarding the estimated spatial coordinates. Additionally, if our task is imputation, meaning that some spatial coordinates may be exactly available, the solution to the estimation is, therefore, an optimization problem with constraints that these exact coordinates must match.
@@
$\min_{\widehat{s}_{i}}\quad$ $\sum_{i,j}\left|\|z_{i}-z_{j}\|-\|\widehat{s}_{i}-\widehat{s}_{j}\|\right|$ (6)
@@
The objective function (6) is highly similar to the distance-preserving loss defined in (3). Observe (3) encodes distance-preserving property into the learned latent embedding, while (6) extracts the embedded distance-preserving property to reconstruct spatial information. This can be viewed as an implicit second set of encode-decode-structure, which we specifically utilize, aside from the first set of encoder-decoder-structure in VAE. However, a subtle difference between the two losses is that the parameters to be optimized in (6) and (3) are fundamentally different, where (6) is parameterized by the spatial information $s_{i}$, and (3) is parametrized by the parameters in the encoder network $\theta$.
@@
The cardinality of the observed spatial information index set $\mathcal{I}$ plays a critical role in the solution of the optimization problem. The optimal solution to (6) is only unique when $\left|\mathcal{I}\right|\geq 3$ but otherwise not, as outlined in Figure 4. When $\left|\mathcal{I}\right|=0$, the problem boils down to the classical multidimensional scaling (MDS) [41] objective, which is a well-studied dimensionality reduction problem and can be solved via, e.g., the SMACOF algorithm [27].
Additionally, when $\left|\mathcal{I}\right|=0$, the optimization procedure can also be skipped: By directly setting the dimension of the latent space to $d=2$, the learned latent embedding can then be naturally regarded as the estimated spatial information, i.e. $z_{i}=\widehat{s}_{i}$. However, when the extracted latent representation is intended for other uses, directly setting $d=2$ is not recommended. This is because the dimension of the latent embedding is physically interpreted as the intrinsic dimension of the manifold where the gene expression features lie, which is typically greater than two. Deliberately setting it to two may result in over-compression of the gene expression feature, resulting in information loss and making the learned embeddings ineffective for other uses. Tuning for the latent dimension could also require a changed set of training hyperparameters, which should be cautious.
The whole algorithm is summarized in Algorithm 1. In the ideal case where the algorithm is capable of fully capturing the distribution, the recover the real spatial information matrix invariant under translation, scaling, and rotation. This invariance is sufficient in practice because the relative spatial relationships between cells are preserved, which is the key information needed for most spatial transcriptomics analyses. The absolute coordinates are often arbitrary and less important than the relative positions and distances between cells in the tissue context. We will describe how such similarity can be evaluated in the experiment section.

### v3

# Distance-Preserving Representations for Genomic Spatial Reconstruction

@@ Abstract
Representation learning, spatial transcriptomics, single-cell RNA sequences, variational inference, and dimensionality reduction.
Figure 1: The model architecture of dp-VAE with vanilla VAE as its backbone. During the training stage, the distance-preserving loss ($\mathcal{L}_{\textrm{DP}}$) enforces distance preservation between the spatial domain ($S$) and the latent space ($Z$), enforcing the encoder network (e.g., parameterized as a fully connected neural network) to capture meaningful signals of the tissue spatial characteristics of the provided gene expression ($X$). During the inference/test stage, the extracted latent representations are fed to an optimizer, such as multidimensional scaling (MDS), if it is a spatial reconstruction task instead of a spatial context imputation task.
@@ 1 Introduction
Understanding the tissue spatial context from which the gene expressions are extracted is crucial to many applications, deepening the understanding of cellular organizations and their interactions within tissues, as well as illuminating biological insights in neuroscience, developmental biology, precise diagnostics, personalized therapies, and a range of disease mechanisms such as cancer [yu2022spatial]. As a result, measurement technologies such as spatial transcriptomics have emerged as a powerful tool in modern biology, bridging the gap between single-cell genomics and the larger physiological processes at play in living tissues.
However, various practical and technical constraints limit the range of spatial context and the granularity that these measurement tools can directly capture. For example, spatial transcriptomics (ST) technologies such as Visium by 10x Genomics and NanoString GeoMx [dries2021advances] typically do not have single-cell resolution [chen2023spatial], making the extracted spatial coordinates within the tissue space too coarse for many precision-focused downstream analyses. Enhancing resolution to this degree requires specialized equipment, intricate protocols, and substantial computational infrastructure, thereby limiting its availability to highly specialized institutions and facilities. On the other hand, the usual single-cell RNA sequencing (scRNA-seq) is capable of extracting gene expression at a localized single-cell level but lacks spatial contexts. Given observations of the complementary nature of the pros and cons of these two existing mainstream gene expression extraction technologies, the following question can be asked:
@@
An approach towards solving this question can potentially avoid the need for the deployment of spatial transcriptomics, making the tasks we intend to carry out with spatial context more accessible.
As an initial step toward tackling this question, we propose a generic representation learning and transfer learning framework enabling the inference of spatial context from purely gene expression data (e.g., collected from scRNA-seq) by drawing reference from datasets where gene expression is paired with spatial information (e.g., collected from spatial transcriptomics). This procedure requires little or even no reliance on ST technologies, depending on the type of reference data that the user intends to use, which can even come from existing publicly available databases. Specifically, our representation learning framework is based on the variational autoencoder (VAE) models [vae], consisting of an encoder and a decoder network and trained on the reference datasets. Instead of using the standard evidence lower bound (ELBO) loss, we propose to regularize it with a distance-preserving loss function, which utilizes the spatial context as auxiliary information to enforce the learned representation to be geometrically similar to the reference dataset. This regularization helps the model capture useful marker genes and signals from gene expressions that are representative of their spatial contexts. During inference, since spatial context is not used as the input to the model, the model can be applied to gene expression datasets without spatial context information, but may share similar expression and geometric patterns with the reference dataset, thereby effectively extrapolating to the task of spatial reconstruction and imputation of the missing spatial context in these inference datasets. Next, the spatial reconstruction task can be formulated as an optimization that reversely solves for the optimal coordinate sets aligning with the distance geometry of extracted latent representations from the input gene expression data. This optimization can be bypassed when we set the latent representation space to match the dimensionality of the tissue’s spatial domain. An illustration of the pipeline is provided in Figure 1.
@@ 2 Related Works
Spatial transcriptomics analysis has consistently garnered interest from the statistical machine learning community. It includes different tasks such as cellular deconvolution, clustering, visualization, cell classification, batch correction, gene imputation, and differential expression [ma2022spatially, long2023spatially, lopez2018deep, du2020model]. Classical statistical models such as marked point process [edsgard2018identification], Gaussian process [svensson2018spatialde], generalized linear spatial models [sun2020statistical, zhu2021spark], and autoregression [ma2022spatially] have been deployed for these modeling tasks, where spatial context information is typically integrated into these methods in the form of covariance matrices. In the meantime, machine learning models empowered with deep neural networks can instead directly leverage the spatial context information. For example, convolution neural network-based methods such as CoSTA [xu2021costa], and/or graph-based methods using graph neural networks such as GLISS [zhu2020integrative], SpaGCN [hu2021spagcn], GraphST [long2023spatially]. See [zeng2022statistical] for a survey of these methods.
Our work aligns more closely with literature that explores how various generative models (e.g. autoencoders) can be applied to learning efficient representation from ST and scRNA-seq, such as STAGATE [dong2022deciphering], scVSC [wang2024scvsc], scVI [lopez2018deep], and gimVI [lopez2019joint]. Some studies have also proposed spatial reconstruction methodologies of tissue spatial structure, such as Tangram [biancalani2021deep], and NovoSpaRc [moriel2021novosparc]. However, to the best of our knowledge, our work is the first to propose to bridge representation learning with tissue spatial reconstruction, drawing reference from out-of-sample and out-of-distribution datasets. Additionally, the previously mentioned statistical and machine learning models dedicated to spatial transcriptomics are also inapplicable to such tasks, as the spatial contexts are required as a conditional input to the models during the inference stage.
The distance-preserving loss used in our study has (concurrently) emerged in parallel domains within the broader machine-learning literature. For instance, distance-preserving mappings have been introduced for GANs for unsupervised domain mapping [benaim2017one], image-to-audio transformation [port2020earballs, kim2020face]. It has also appeared as a regularizer for VAEs [chen2022local] and adversarial learning [hadsell2006dimensionality, mao2019metric]. A stronger notation compared to distance-preserving, named isometry [yonghyeon2021regularized, lee2021neighborhood, beshkov2022isometric] has also been extensively studied in machine learning models such as VAEs. Building on top of these works, our derivation of the distance-preserving loss’s connection to distortion or the bi-Lipschitz condition contributes to formalizing this notion.
@@ 3 Problem setup
At the high level, the goal of representation learning is to learn a mapping that extracts latent embeddings $z\in\mathbb{R}^{d}$ from given gene expression $x\in\mathbb{R}^{D}$. In our setting, we additionally assume that we have access to some spatial context information $s\in\mathbb{R}^{2}$ associated with each gene expression, indicating where the cell is located in the tissue slice. It is only available in the training stage, but not during inference.
We expect that the learned representation learning model can also undergo some procedure to reconstruct the spatial domain given only gene expression during the inference stage, which we refer to as the spatial reconstruction task. Formally, given the gene expression a set of gene expressions denoted as $x_{i}\in\mathbb{R}^{D}$, $i=1,\ldots,N$, the goal of spatial reconstruction is to obtain a mapping to infer the spatial coordinates $s_{i}\in\mathbb{R}^{2}$ for each cellular location $i$. In some cases, the spatial information of some cell has already been observed, indexed by $i\in\mathcal{I}\subseteq\left\{1,\ldots,N\right\}$, then the spatial reconstruction task turns into the spatial imputation task, which aims to infer the rest of the spatial coordinates $s_{i}$, where $i\in\left\{1,\ldots,N\right\}/\mathcal{I}$. The model should be able to handle both spatial reconstruction and imputation tasks. Given their similarity, we use the two terms interchangeably, referring to them as spatial reconstruction in most cases throughout this paper.
We remark that, in practice, both spatial reconstruction and imputation are useful for many downstream analyses, such as spatial domain identification, region-region interaction analysis, and spatial trajectory analysis. The inference task is similar in spirit to the setting of transfer learning [pan2009survey].
@@ 4 Method
To tackle the aforementioned challenges, we propose a gene expression representation learning framework. The model, named distance-preserving variation autoencoder (dp-VAE), incorporates the spatial information into the learning objective as a penalty for the error of the pairwise similarity between the model’s learned latent variable and the spatial information of the training data. In this case, the spatial information can also be viewed as an auxiliary variable during the training stage rather than a direct input of the model, therefore it is not required during the inference stage [^1] Whereas conditional models such as CVAE [sohn2015learning, wu2023counterfactual] do require the spatial context as inputs, therefore cannot be applied in this setting.. During inference, spatial reconstruction, and spatial imputation, this penalty can be used reversely to decipher the spatial coordinates from the extracted latent representations of the data.
@@ 4.1 Background: Variational Autoencoder
However, the existing encoding paradigm presented in (2) reveals implicit biases, undermining its efficacy. Two primary concerns highlight this limitation: ($i$) The likelihood-focused optimization objective, while capturing some meaningful signals, underscores probabilistic aspects that could be incorrectly specified due to cross-dependencies or non-identical distributed genomic samples. This could lead to entangled and biased representations of spatial context being learned from the training data; ($ii$) In the meantime, directly inserting the spatial information as part of the input variable violates our assumption about the setting where spatial context is not available as input during the inference stage, but rather should be the output of the model. Therefore, there is a need to develop a new method that incorporates likelihood-based generative modeling strategies with spatial information.
@@ 4.2 Distance-Preserving Variational Autoencoder
where $\left\|\cdot\right\|$ is the Euclidean norm. The intuition behind the design of this loss is to ensure that the latent space preserves the relative pairwise distances between data points in the original space, while the scaling factor $\lambda$ can be adjusted appropriately to account for differences in the variance of the underlying prior distributions of the latent and original spaces. This encourages the model to maintain geometric consistency while learning meaningful representations of gene expression aligned with the objective of VAE. Next, we incorporate this loss as a weighted regularizer in the objective function of the training model
@@
where the reconstruction term $\mathcal{L}_{\rm recon}$ and KL divergence term $\mathcal{L}_{\rm KL}$ are as defined in (2). The outer-level regularization coefficients $\alpha_{1},\alpha_{2}\leq 0$ control for the numerical balance of the distance-preserving loss and the ELBO loss, and $\beta$ is the temperature coefficient for the KL-divergence term [higgins2017beta]. The model is trained by learning the parameters of the encoder and decoder networks $\theta$ and $\phi$ as well as the scaling parameter $\lambda$ to minimize (4).
@@
    From the biological perspective, neighboring cells, often from similar tissue types, tend to exhibit homogeneity in attributes like gene expression and biological functions. Representation-based generative models, which aim to capture such spatial relationships and tissue-level patterns, should utilize such geometric structures in learning disentangled representations that are meaningful. Prior works [hu2021spagcn, sun2020statistical] have illustrated that proximity in physical distance between cells has a strong association with their gene expressivity. This is also the initial of many spatial transcriptomics machine learning models to incorporate spatial context as inputs to enhance the modeling of cellular behavior and tissue organization, enabling more robust, context-aware insights into biological systems. If viewed reversely, it is also reasonable to claim that the learned gene expression representations can also reveal the associated spatial context from where they are located in the tissue.
  • (iiii)
@@
It is noted that the first point is also supported by a simple experiment we conducted, illustrated in Figure 2. It can be seen that out of the 27 datasets that we are using (see the experiment section for a detailed description), 23 datasets show a statistically significant ($95\%$) positive correlation between the spatial distance of the sample locations and the gene expression similarity. This highlights the potential of seeing the spatial context as a useful source of information to aid the learning of gene expression representations, or, conversely, gene expression representations are capable of revealing information about their associated spatial context. Though we want to note that, despite the statistical significance shown in Figure 2, the magnitude of the correlations differs across datasets, with some close to zero. This means that the theoretical maximum amount of information that can be extracted from the association could be weak for certain datasets, which could lead to high variance and instability for poorly constructed estimation models. This further motivates our study, which aims at constructing a powerful machine learning architecture that exploits such information for accurate spatial reconstruction and imputation.
@@
[…]
@@ 4.3 Spatial Reconstruction and Imputation
In the next step, we aim to carry out the spatial reconstruction and imputation task, where we only have inference data consisting of gene expression but not spatial context information. Assuming that we have access to some gene expression data that are paired with spatial context information and are statistically similar to the inference data, then dp-VAE can be trained on a reference dataset, but transferred to the inference data for application. Specifically, since dp-VAE encodes the spatial context patterns in the latent representation via a distance-preserving regulated mapping, we can reversely find the optimal spatial coordinates in the spatial domain that align with the distance patterns in the extracted latent space. This can be formulated as an optimization problem regarding the estimated spatial coordinates. Additionally, if our task is imputation, meaning that some spatial coordinates may be exactly available, the solution to the estimation is, therefore, an optimization problem with constraints that these exact coordinates must match.
@@
[…]
@@
The objective function (6) is highly similar to the distance-preserving loss defined in (3). Observe (3) encodes the distance-preserving property into the learned latent embedding, while (6) extracts the embedded distance-preserving property to reconstruct spatial information. This can be viewed as an implicit second set of encoder-decoder-structure, which we specifically utilize, aside from the first set of encoder-decoder-structure in VAE. However, a subtle difference between the two losses is that the parameters to be optimized in (6) and (3) are fundamentally different, where (6) is parameterized by the spatial information $s_{i}$, and (3) is parametrized by the parameters in the encoder network $\theta$.
@@
The cardinality of the observed spatial information index set $\mathcal{I}$ plays a critical role in the solution of the optimization problem. The optimal solution to (6) is only unique when $\left|\mathcal{I}\right|\geq 3$ but otherwise not, as outlined in Figure 4. When $\left|\mathcal{I}\right|=0$, the problem boils down to the classical multidimensional scaling (MDS) [saeed2018survey] objective, which is a well-studied dimensionality reduction problem and can be solved via, e.g., the SMACOF algorithm [leeuw1977application].
Additionally, when $\left|\mathcal{I}\right|=0$, the optimization procedure can also be skipped: By directly setting the latent space dimension to $d=2$, the learned embedding naturally represents estimated spatial coordinates ($z_{i}=\widehat{s}_{i}$), and could be naturally adopted in spatial reconstruction and imputation tasks. However, when latent representations are intended for other applications, such as domain identification, trajectory inference, and gene expression imputation, $d=2$ is not recommended, as the intrinsic dimension of gene expression manifolds typically exceeds two. Deliberately constraining to two dimensions may over-compress features, causing information loss and reducing effectiveness for these tasks. Tuning for the latent dimension could also require a changed set of training hyperparameters, which should be handled with caution.
The whole algorithm is summarized in Algorithm 1. In the ideal case where the algorithm is capable of fully capturing the distribution, it recovers the real spatial information matrix invariant under translation, scaling, and rotation. This invariance is sufficient in practice because the relative spatial relationships between cells are preserved, which is the key information needed for most spatial transcriptomics analyses. The absolute coordinates are often arbitrary and less important than the relative positions and distances between cells in the tissue context. We will describe how such similarity can be evaluated in the experiment section.